Ricky drove from Town A to Town B in 3 hours. His return trip from Town B to Town A took 5 hours because he drove 15 miles per hour slower on the return trip. How fast did Ricky drive on the trip from Town A to Town B?
Please explain step by step. I am having a really hard time grasping distance problems. I know d=rt. I was thinking something like 3x and 5x-15 but I'm just not sure how to set it up. PLEASE help.

Answer by richwmiller(17219)

3x is good
The return speed is x-15 so
5*(x-15)
The distance is the same in both directions
Set them equal
3x=5*(x-15)
We want solve for x the speed from A to B
x=75/2=37.5